# Reconstruction of frequency-localized functions from pointwise samples via least squares and deep learning

### v1

@@ 4 Numerical experiments
We assess the performance of Least Squares and Deep Learning for the reconstruction of the one-dimensional function
$f_{1}(x):=\cos(10x)e^{-\pi x^{2}},\qquad\forall x\in[-1,1],$
and the two-dimensional function
@@
The one-dimensional function $f_{1}$, defined as the product of a trigonometric function and a Gaussian, was considered in [28] as a simple example in which the Slepian basis achieves better approximation accuracy than the standard Fourier basis. The two-dimensional function $f_{2}$ generalizes the one-dimensional function $f_{1}$ while reducing the cosine frequency from 10 to 0.2. Both functions are frequency-localized since most of their energy is concentrated within a compact interval in the frequency domain.
Training set, test set, and error metric.
For a given target function $f$ that we seek to approximate using either Least Squares or Deep Learning, we generate data sets of the form $\{(x_{j},f(x_{j})\}_{j=1}^{m}$ with $x_{j}$ drawn independently from the uniform distribution on $[-1,1]^{d}$ and the sample size $m$ taking values up to $10000$. For a set of $m$ training points, we also consider a test set $\{x_{j}^{\text{test}}\}_{j=1}^{m_{\text{test}}}$ of size $m_{\text{test}}=0.2m$, sampled independently according to the uniform distribution on $[-1,1]^{d}$. The corresponding test error is given by the Root Mean Square Error (RMSE), defined as
where $f$ is the target function, and $\tilde{f}$ is the approximation obtained from either method. Specifically, adopting the notation in Section 3, $\tilde{f}=f^{\natural}$ for Least Squares, and $\tilde{f}={\rm R}(\Psi^{\natural})$ for Deep Learning, where $\Psi^{\natural}$ is the trained NN.
Slepian basis functions.
To construct one-dimensional Slepian basis functions $\varphi_{\mathsf{w},j}$, we employ the Python package scipy.special [45], which is based on the angular solutions of the Helmholtz wave equation of the first kind. In fact, each function $\varphi_{\mathsf{w},j}$ is a scalar multiple of a corresponding angular solution [28]. To construct two-dimensional Slepian basis functions, we utilize the tensor-product definition (2.7) $\varphi_{\mathsf{w},\vec{\nu}}=\varphi_{\mathsf{w},\nu_{1}}\otimes\varphi_{\mathsf{w},\nu_{2}}$. Finally, in both cases, we take the index set to be the hyperbolic cross $\Lambda^{\text{HC}}_{n-1}$ (Definition 2.2) for various orders $n\in\mathbb{N}$.
@@
Following standard deep learning practice, we consider fully trained NNs. Here, we deviate from the theoretical setting of Theorem 3.2, where only the last layer is trained. NN architectures are built according to the ratio $r=L/N=0.1$, where $L$ represents the number of hidden layers and $N$ the number of neurons per layer. Specifically, we conduct experiments for $L=1,\dots,10$, with $r=0.1$, a choice empirically supported by [4]. Weights and biases are initialized from a normal distribution with mean 00 and standard deviation $0.1$. Training is performed using Adam [24] with an exponential decay rate of $0.85$ for $150$ epochs. To ensure the robustness and reliability of the results, each experiment is repeated $20$ times.
@@
Figure 2 illustrates the results of our first experiment on approximating the one-dimensional function $f_{1}$. In accordance with our theoretical guarantees, both methods successfully approximate this frequency-localized function, with the RMSE converging as the number of sampling points increases. Moreover, the number of parameters in each reconstruction model significantly affects the RMSE. For Least Squares, increasing the number of Slepian basis functions (corresponding to increasing the order $n$) generally lowers test errors, with the only recorded exception being $n=36$. Similarly, for Deep Learning, increasing the number of network layers $L$ results in a noticeable error reduction. However, note that Least Squares typically requires fewer samples as well as significantly fewer parameters to achieve high accuracy. The best accuracy achieved by Least Squares (RMSE $\approx 10^{-12}$) is several orders of magnitude higher than that by Deep Learning (RMSE $\approx 10^{-4}$). In summary, in one dimension, Least Squares is much more effective than Deep Learning for the approximation of $f_{1}$.
Figure 2: Comparison of reconstruction methods for $f_{1}(x)=\cos(10x)e^{-\pi x^{2}}$ from samples. The left panel shows results using Deep Learning, while the right panel shows results using Least Squares. Here, $L$ denotes the number of layers in each NN, and $n$ denotes the order of the hyperbolic cross $\Lambda^{\rm HC}_{n-1}$.
@@
Figure 3 illustrates the results of our second experiment on approximating the two-dimensional function $f_{2}$. Both Least Squares and Deep Learning successfully approximate this frequency-localized function, in line with our theory. As in the one-dimensional case, we observe RMSE convergence as a function of the number of samples $m$. However, there are notable differences between the one- and two-dimensional cases. First, the best accuracy achieved by both methods is now much more similar. Both Least Squares and Deep Learning achieve a best mean RMSE of around $10^{-4}$. Interestingly, in some experiments, Deep Learning can attain a lower RMSE, as indicated by the size of the variance regions (the shaded regions in Figure 3, left). Hence, the accuracy gap between Least Squares and Deep Learning is much less pronounced in dimension two than in dimension one. Second, although the advantage of adding parameters to the model (for both Least Squares and Deep Learning) is still there, we observe some intriguing numerical phenomena. For Least Squares, considering a larger Slepian basis (i.e., larger order $n$) might not lead to a better accuracy (see the curve corresponding to $n=26$). Meanwhile, for Deep Learning, the benefit of increasing $L$ for accuracy becomes limited after $L\geq 6$. In summary, dimensionality appears to play a crucial role when comparing Least Squares and Deep Learning. While a comprehensive numerical study is beyond the scope of this paper, our preliminary results suggest that the performance difference between the two reconstruction methods could increase in dimension three and higher.
Figure 3: Comparison of reconstruction methods for $f_{2}(x)=\cos(0.2x)\cos(0.2y)e^{-\pi(x^{2}+y^{2})}$ from samples. The left panel shows results using Deep Learning, while the right panel shows results using Least Squares. Here, $L$ denotes the number of layers in each NN, and $n$ denotes the order of the hyperbolic cross $\Lambda^{\rm HC}_{n-1}$.
@@ 5 Proofs
To facilitate the proofs of Theorem 3.1 and Theorem 3.2, detailed in Subsection 5.2 and Subsection 5.4 respectively, we first establish several key preliminary results in Subsection 5.1 and Subsection 5.3. Both theorems necessitate an understanding of the uniform norm of a normalized Slepian basis function within the centered $d$-dimensional unit cube. Additionally, Theorem 3.2 relies on the polynomial approximability of neural networks and fundamental network operations, which are outlined in Appendix A for clarity and reference.
@@ 5.1 A preliminary result for the main theorems
In what follows, we recall from (3.1) the simplified notation $\gamma(\mathsf{w})=\lceil\log_{2}(30\mathsf{w})\rceil$.
@@
see for example [38, Remark 2.4]. However, (5.2) holds only asymptotically (and a rectified version of (5.2) is given in [48, Theorem 2.7]). Specifically, for sufficiently large values of $j$, beyond the critical index $j^{\star}(\mathsf{w})$, the global maximum of $|\varphi_{\mathsf{w},j}|$ on $[-1,1]$ shifts to the boundary of the interval, becoming equal to $|\varphi_{\mathsf{w},j}(1)|$, which is of order $\mathcal{O}(\sqrt{j})$, as numerically demonstrated in [48, Theorem 8.4]. This shifting behavior, as discussed in both [48, Theorem 3.38], [8, Theorem 3.1], occurs when $\mathsf{w}^{2}/\chi_{\mathsf{w},j}\leq 1$. The second theorem will be later utilized in our proof of Proposition 5.2. For low indices $j$, [8, Remark 3.1] implies that $\|\varphi_{\mathsf{w},j}\|_{L^{\infty}([-1,1])}\leq C\sqrt{\mathsf{w}}$. Notably, $\varphi_{\mathsf{w},0}$ is even, positive on the interval $(-1,1)$ and attains its global maximum at the origin. Yet, for low indices $j$, including $j=0$, the behavior of $\varphi_{\mathsf{w},j}(0)$ is poorly understood.
@@
$h(x):=\frac{\log\big{(}30\mathsf{w}\mathbbm{1}_{\{x\leq j^{\star}(\mathsf{w})\}}+\tilde{c}(x+1)\mathbbm{1}_{\{x>j^{\star}(\mathsf{w})\}}\big{)}}{\log(x+1)}.$ (5.4)
@@
$h(x)$ is nonincreasing on $[1,\infty)$.
Observe that the claim is evident for either the case $1\leq x\leq j^{\star}(\mathsf{w})$ or for the case $x>j^{\star}(\mathsf{w})$. Indeed, by definition (5.4), on $[1,j^{\star}(\mathsf{w})]$,
$h(x)=\frac{\log(30\mathsf{w})}{\log(x+1)}$
is decreasing, and on $(j^{\star}(\mathsf{w}),\infty)$,
$h(x)=\frac{\log(\tilde{c}\,(x+1))}{\log(x+1)}=\frac{\log(\tilde{c})}{\log(x+1)}+1$
is also decreasing. Hence it suffices to show, if $11\leq x_{1}\leq j^{\star}(\mathsf{w})<x_{2}\leq j^{\star}(\mathsf{w})+$ 1, then $h(x_{2})\leq h(x_{1})$. Write $x_{2}=j^{\star}(\mathsf{w})+\varepsilon=\lfloor 4\mathsf{w}\rfloor-1+\varepsilon$ for some $\varepsilon\in(0,1]$. Given that $\mathsf{w}\geq 1$, and so $30\mathsf{w}\geq 6(4\mathsf{w}+1)\geq\tilde{c}(j^{\star}(\mathsf{w})+\varepsilon+1)$, we gather
$h(x_{1})=\frac{\log(30\mathsf{w})}{\log(x_{1}+1)}\geq\frac{\log(\tilde{c}(j^{\star}(\mathsf{w})+\varepsilon+1))}{\log(j^{\star}(\mathsf{w})+\varepsilon+1)}=h(x_{2}),$
thereby affirming the validity of Claim 5.4. As a consequence, for all $x\geq 1$,
$\frac{\log\big{(}30\mathsf{w}\mathbbm{1}_{\{x\leq j^{\star}(\mathsf{w})\}}+\tilde{c}(x+1)\mathbbm{1}_{\{x>j^{\star}(\mathsf{w})\}}\big{)}}{\log(x+1)}\leq\frac{\log(30\mathsf{w})}{\log(2)}=\gamma(\mathsf{w})$
or equivalently,
$30\mathsf{w}\mathbbm{1}_{\{1\leq x\leq j^{\star}(\mathsf{w})\}}+\tilde{c}(x+1)\mathbbm{1}_{\{x>j^{\star}(\mathsf{w})\}}\leq(x+1)^{\gamma(\mathsf{w})}.$ (5.5)
Combining (5.5), (5.3), we can infer that
@@
To prove Proposition 5.2, we rely on the following two well-known lemmas. The first, Lemma 5.5 (see, for instance, [26, 27]) concerns the spectral gap in the eigenvalues $\mu_{\mathsf{w},j}$ (2.4), while the second, Lemma 5.6, pertains to the magnitude of $\varphi_{\mathsf{w},j}$ for sufficiently large index $j$. The proof of Proposition 5.2 will come after the presentation of these lemmas.
@@
Recall that the Slepian basis functions $\varphi_{\mathsf{w},j}$ serve as eigenfunctions for the operator $\mathcal{L}_{\mathsf{w}}$ (2.5), with the corresponding eigenvalues denoted by $\chi_{\mathsf{w},j}$ (2.6). Let $j\geq\frac{2\mathsf{w}}{\pi}$. Then it follows from [31, Theorem 4.4] that
@@
On the one hand, when (5.8) is satisfied, we can employ [31, Theorem 3.39] to obtain
@@
On the other hand, according to [8, Theorem 3.1], (5.8) also yields
@@
Extend the uniform measure ${\bf u}$ on $[-1,1]$ to a measure on $\mathbb{R}$ by setting ${\rm d}{\bf u}(t)=\frac{1}{2}{\rm d}t$. We show that $\{\varphi_{\mathsf{w},j}\}_{j\in\mathbb{N}_{0}}$ forms an orthogonal system in $L_{\bf u}^{2}(\mathbb{R})$. Since $\varphi_{\mathsf{w},j}$ is an eigenfunction of the operator $\mathcal{Q}_{\mathsf{w}}$ (2.3) corresponding to the eigenvalue $\mu_{\mathsf{w},j}$,
$\mathcal{Q}_{\mathsf{w}}\varphi_{\mathsf{w},j}(x)=\int_{-1}^{1}\varphi_{\mathsf{w},j}(t)\,\frac{\sin\mathsf{w}(x-t)}{\pi(x-t)}\,{\rm d}t=\mu_{\mathsf{w},j}\varphi_{\mathsf{w},j}(x),\qquad\forall x\in\mathbb{R}.$ (5.11)
Hence, we deduce that
$\int_{\mathbb{R}}\varphi_{\mathsf{w},j}(x)\varphi_{\mathsf{w},k}(x)\,{\rm d}{\bf u}(x)$
$=\frac{1}{2\mu_{\mathsf{w},j}\mu_{\mathsf{w},k}}\int_{[-1,1]\times[-1,1]}\varphi_{\mathsf{w},j}(t)\varphi_{\mathsf{w},k}(u)\bigg{(}\int_{\mathbb{R}}\frac{\sin\mathsf{w}(x-t)}{\pi(x-t)}\,\frac{\sin\mathsf{w}(x-u)}{\pi(x-u)}\,{\rm d}x\bigg{)}\,{\rm d}t{\rm d}u$
$=\frac{1}{2\mu_{\mathsf{w},j}\mu_{\mathsf{w},k}}\int_{[-1,1]\times[-1,1]}\varphi_{\mathsf{w},j}(t)\varphi_{\mathsf{w},k}(u)\frac{\sin\mathsf{w}(t-u)}{\pi(t-u)}\,{\rm d}t{\rm d}u$
$=\frac{1}{\mu_{\mathsf{w},j}}\int_{[-1,1]}\varphi_{\mathsf{w},j}(t)\varphi_{\mathsf{w},k}(t)\,{\rm d}{\bf u}(t).$ (5.12)
Above, we have applied a reproducing kernel trick [7, Subsection 3.1.2] at the second inequality and (5.11) at the first and last. Due to the orthonormality of $\{\varphi_{\mathsf{w},j}\}_{j\in\mathbb{N}_{0}}$ in $L_{\bf u}^{2}([-1,1])$, (5.1) evaluates to zero if $j\neq k$ and to $\frac{1}{\mu_{\mathsf{w},j}}$ if $j=k$, we conclude that $\{\varphi_{\mathsf{w},j}\}_{j\in\mathbb{N}_{0}}$ indeed constitutes an orthogonal system in $L_{\bf u}^{2}(\mathbb{R})$ [^1] In other words, the Slepian functions are bi-orthogonal in the sense that they are orthogonal over both a given finite interval and $\mathbb{R}$.. In particular,
$\|\varphi_{\mathsf{w},j}\|_{L_{\bf u}^{2}(\mathbb{R})}=\sqrt{\frac{1}{\mu_{\mathsf{w},j}}}.$ (5.13)
Hence, by applying Lemma 5.5 to (5.13), we obtain for $j\leq\lfloor 4\mathsf{w}\rfloor-1$ that
$\|\varphi_{\mathsf{w},j}\|_{L_{\bf u}^{2}(\mathbb{R})}\leq\sqrt{2}.$ (5.14)
@@
[…]
and subsequently, for $j\in\mathbb{N}_{0}$,
$\|\varphi_{\mathsf{w},j}\|_{L^{\infty}(\mathbb{R})}\leq\sqrt{\frac{2\mathsf{w}}{\pi}}\|\varphi_{\mathsf{w},j}\|_{L_{\bf u}^{2}(\mathbb{R})}.$ (5.15)
Then combining (5.14) with (5.15) for $j\leq\lfloor 4\mathsf{w}\rfloor-1$ yields
@@
$\alpha=\alpha(\mathcal{S}_{\Lambda},\{\vec{y}_{j}\}_{j=1}^{m})\\ :=\bigg{\{}\frac{1}{m}\sum_{j=1}^{m}|g(x_{j})|^{2}:g\in\mathcal{S}_{\Lambda}\text{ and }\|g\|_{L_{{\bf u}}^{2}([-1,1]^{d})}=1\bigg{\}}.$
To simplify the notation, we let $K:=\texttt{\#}\Lambda$ and assign a fixed order to the elements of $\Lambda$ via a bijection $\Pi:\{1,\dots,K\}\to\Lambda$. We rewrite $\varphi_{\mathsf{w},\vec{\nu}}$ as $\varphi_{\mathsf{w},\Pi(k)}$ temporarily, using $k=1,2,\dots,K$ as indices. Then, since $\mathcal{S}_{\Lambda}$ is a finite-dimensional space, with a basis formed by the vectors $\{\varphi_{\mathsf{w},\Pi(k)}\}_{k=1}^{K}$, we apply the Courant-Fisher min-max theorem [22, Theorem 4.2.6] to get
@@
where $A\in\mathbb{R}^{K\times m}$ is given by
$A_{jk}:=\dfrac{1}{\sqrt{m}}\varphi_{\mathsf{w},\Pi(k)}(\vec{y}_{j}).$ (5.16)
Here, $\sigma_{\rm min}(A)$ denotes the minimum singular value of $A$ and $\lambda_{\rm min}(A^{*}A)$ the minimum eigenvalue of $A^{*}A$. A positive lower bound on $\alpha$, crucial for the existence of a unique solution to (2.8), can be further guaranteed via Monte Carlo sampling if $m$ is sufficiently large. This is formalized in the following theorem, adapted from [3, Theorem 5.7].
@@
Let $\beta,\delta\in(0,1)$ and set [^2] The quantity $\kappa(\mathcal{S}_{\Lambda})$ in (5.17) is known as the $L^{\infty}$-norm of the Christoffel function of $\mathcal{S}_{\Lambda}$ [3, Definition 5.4].
$\kappa(\mathcal{S}_{\Lambda}):=\max_{\vec{y}\in[-1,1]^{d}}\sum_{k=1,\dots,K}|\varphi_{\mathsf{w},\Pi(k)}(\vec{y})|^{2}.$ (5.17)
If $m\geq c_{\delta}\kappa(\mathcal{S}_{\Lambda})\log(K/\beta)$, then with probability at least $1-\beta$, the discrete stability constant $\alpha=\alpha(\mathcal{S}_{\Lambda},\{\vec{y}_{j}\}_{j=1}^{m})$ satisfies
@@
Assuming the validity of (5.18), the theory outlined in [3, Section 5], particularly [3, Corollary 5.9], demonstrates that every $f\in\mathcal{C}([-1,1]^{d})$ has a unique least squares approximation $f^{\natural}$ in (2.8) that satisfies
$<\Big{(}1+\frac{1}{\sqrt{1-\delta}}\Big{)}\inf_{g\in\mathcal{S}_{\Lambda}}\|f-g\|_{L^{\infty}([-1,1]^{d})}+\frac{1}{\sqrt{1-\delta}}\|\vec{e}\|_{2}.$ (5.19)
It then follows from Theorem 5.7 that the estimation of $\kappa(\mathcal{S}_{\Lambda})$ is the key piece to guarantee (5.2) with probability at least $1-\beta$. Evidently from (5.17), the size of $\kappa(\mathcal{S}_{\Lambda})$ is impacted by the choice of $\Lambda$. When $\Lambda$ is a hyperbolic cross in low dimensions, an upper bound for $\kappa(\mathcal{S}_{\Lambda})$ can be provided by the next proposition.

### v2

@@ 4.2 Comparisons of different initializations for Deep Learning
We assess the performance of Least Squares and Deep Learning for the reconstruction of the one-dimensional function $f_{1}$ defined in (4.1) and the two-dimensional function
@@
The two-dimensional function $f_{2}$ generalizes the one-dimensional function $f_{1}$ while reducing the cosine frequency from 10 to 0.2. Both functions are frequency-localized since most of their energy is concentrated within a compact interval in the frequency domain.
For each function $f$, we consider the sample size $m$ taking values up to $10000$. For a set of $m$ training points, we also consider a test set $\{x_{j}^{\text{test}}\}_{j=1}^{m_{\text{test}}}$ of size $m_{\text{test}}=0.2m$, sampled independently according to the uniform distribution on $1,1]^{d}$.
@@
Following standard deep learning practice, we consider fully trained NNs. Here, we deviate from the theoretical setting of Theorem 3.3, where only the last layer is trained. NN architectures are built according to the ratio $r=L/N=0.1$, where $L$ represents the number of hidden layers and $N$ the number of neurons per layer. Specifically, we conduct experiments for $L=1,\dots,10$, with $r=0.1$, a choice empirically supported by [adcock2021gap]. Weights and biases are initialized from a normal distribution with mean 0 and standard deviation $0.1$. Training is performed using Adam [diederik2015adam] with an exponential decay rate of $0.85$ for $150$ epochs. To ensure the robustness and reliability of the results, each experiment is repeated $20$ times.
@@
Figure 6 illustrates the results of our first experiment on approximating the one-dimensional function $f_{1}$.
Figure 6: Comparison of reconstruction methods for $f_{1}(x)=\cos(10x)e^{-\pi x^{2}}$ from samples. The left panel shows results using Deep Learning, while the right panel shows results using Least Squares. Here, $L$ denotes the number of layers in each $N$, and $n$ denotes the order of the hyperbolic cross $\Lambda^{\rm HC}_{n-1}$.
In accordance with our theoretical guarantees, both methods successfully approximate this frequency-localized function, with the RMSE converging as the number of sampling points increases. Moreover, the number of parameters in each reconstruction model significantly affects the RMSE. For Least Squares, increasing the number of Slepian basis functions (corresponding to increasing the order $n$) generally lowers test errors, with the only recorded exception being $n=36$. Similarly, for Deep Learning, increasing the number of network layers $L$ results in a noticeable error reduction. However, note that Least Squares typically requires fewer samples as well as significantly fewer parameters to achieve high accuracy. The best accuracy achieved by Least Squares (RMSE $\approx 10^{-12}$) is several orders of magnitude higher than that by Deep Learning (RMSE $\approx 10^{-4}$). In summary, in one dimension, Least Squares is much more effective than Deep Learning for the approximation of $f_{1}$.
@@
Figure 7 illustrates the results of our second experiment on approximating the two-dimensional function $f_{2}$.
Figure 7: Comparison of reconstruction methods for $f_{2}(x)=\cos(0.2x)\cos(0.2y)e^{-\pi(x^{2}+y^{2})}$ from samples. The left panel shows results using Deep Learning, while the right panel shows results using Least Squares. Here, $L$ denotes the number of layers in each $N$, and $n$ denotes the order of the hyperbolic cross $\Lambda^{\rm HC}_{n-1}$.
Both Least Squares and Deep Learning successfully approximate this frequency-localized function, in line with our theory. As in the one-dimensional case, we observe RMSE convergence as a function of the number of samples $m$. However, there are notable differences between the one- and two-dimensional cases. First, the best accuracy achieved by both methods is now much more similar. Both Least Squares and Deep Learning achieve a best mean RMSE of around $10^{-4}$. Interestingly, in some experiments, Deep Learning can attain a lower RMSE, as indicated by the size of the variance regions (the shaded regions in Figure 7, left). Hence, the accuracy gap between Least Squares and Deep Learning is much less pronounced in dimension two than in dimension one. Second, although the advantage of adding parameters to the model (for both Least Squares and Deep Learning) is still there, we observe some intriguing numerical phenomena. For Least Squares, considering a larger Slepian basis (i.e., larger order $n$) might not lead to a better accuracy (see the curve corresponding to $n=26$). Meanwhile, for Deep Learning, the benefit of increasing $L$ for accuracy becomes limited after $L\geq 6$. In summary, dimensionality appears to play a crucial role when comparing Least Squares and Deep Learning. While a comprehensive numerical study is beyond the scope of this paper, our preliminary results suggest that the performance difference between the two reconstruction methods could increase in dimension three and higher.
@@ 5 Proofs
To facilitate the proofs of Theorem 3.1 and Theorem 3.3, detailed in Section 5.3 and Section 5.5 respectively, we first establish several key preliminary results in Section 5.1 and Section 5.4. Both theorems necessitate an understanding of the uniform norm of a normalized Slepian basis function within the centered $d$-dimensional unit cube. Additionally, Theorem 3.3 relies on the polynomial approximability of neural networks and fundamental network operations, which are outlined in Appendix A for clarity and reference.
@@ 5.1 A preliminary result for the main theorems
In what follows, we recall from (3.1) the simplified notation $\gamma(\mathsf{w})=\lceil\log_{2}(24\mathsf{w})\rceil$.
@@
see for example [shkolnisky2006approximation, Remark 2.4]. However, (5.2) holds only asymptotically (and a rectified version of (5.2) is given in [xiao2001prolate, Theorem 2.7]). Specifically, for sufficiently large values of $j$, beyond the critical index $j^{\star}(\mathsf{w})$, the global maximum of $|\varphi_{\mathsf{w},j}|$ on [1,1][-1,1] shifts to the boundary of the interval, becoming equal to $|\varphi_{\mathsf{w},j}(1)|$, which is of order $\mathcal{O}(\sqrt{j})$, as numerically demonstrated in [xiao2001prolate, Theorem 8.4]. This shifting behavior, as discussed in both [xiao2001prolate, Theorem 3.38], [bonami2014uniform, Theorem 3.1], occurs when $\mathsf{w}^{2}/\chi_{\mathsf{w},j}\leq 1$. The second theorem will be later utilized in our proof of Proposition 5.2. For low indices $j$, [bonami2014uniform, Remark 3.1] implies that $\|\varphi_{\mathsf{w},j}\|_{L^{\infty}([-1,1])}\leq C\sqrt{\mathsf{w}}$. Notably, $\varphi_{\mathsf{w},0}$ is even, positive on the interval (1,1)(-1,1) and attains its global maximum at the origin. Yet, for low indices $j$, including $j=0$, the behavior of $\varphi_{\mathsf{w},j}(0)$ is poorly understood.
@@
$h(x):=\frac{\log\big(24\mathsf{w}\mathbbm{1}_{\{x\leq j^{\star}(\mathsf{w})\}}+\tilde{c}(x+1)\mathbbm{1}_{\{x>j^{\star}(\mathsf{w})\}}\big)}{\log(x+1)}.$
@@
$h(x)$ is nonincreasing on $1,\infty)$.
Observe that the claim is evident for either the case $1\leq x\leq j^{\star}(\mathsf{w})$ or for the case $x>j^{\star}(\mathsf{w})$. Hence it suffices to show, if $1\leq x_{1}\leq j^{\star}(\mathsf{w})<x_{2}\leq j^{\star}(\mathsf{w})+1$, then $h(x_{2})\leq h(x_{1})$, or:
$\frac{\log(24\mathsf{w})}{\log(x_{1}+1)}\geq\frac{\log(\tilde{c}(x_{2}+1))}{\log(x_{2}+1)}.$
Since the right-hand side is a continuous and nonincreasing function, it also suffices if we show
$\frac{\log(24\mathsf{w})}{\log(x_{1}+1)}\geq\lim_{x_{2}\uparrow j^{\star}(\mathsf{w})}\frac{\log(\tilde{c}(x_{2}+1))}{\log(x_{2}+1)}=\frac{\log(\tilde{c}(j^{\star}(\mathsf{w})+1))}{\log(j^{\star}(\mathsf{w})+1)}=\frac{\log(\tilde{c}\lfloor 4\mathsf{w}\rfloor)}{\log(\lfloor 4\mathsf{w}\rfloor)}.$ (5.4)
Note that $\log(\lfloor 4\mathsf{w}\rfloor)=\log(j^{\star}(\mathsf{w})+1)\geq\log(x_{1}+1)$, as $j^{\star}(\mathsf{w})\geq x_{1}$. Therefore (5.4) follows, once $\log(24\mathsf{w})\geq\log(\tilde{c}\lfloor 4\mathsf{w}\rfloor)$ holds. However, this is immediate, since
$\log(24\mathsf{w})\geq\log(23.04\mathsf{w})=\log(\tilde{c}4\mathsf{w})\geq\log(\tilde{c}\lfloor 4\mathsf{w}\rfloor).$
Thus, Claim 5.4 holds, and we get:
$\frac{\log\big(24\mathsf{w}\mathbbm{1}_{\{x\leq j^{\star}(\mathsf{w})\}}+\tilde{c}(x+1)\mathbbm{1}_{\{x>j^{\star}(\mathsf{w})\}}\big)}{\log(x+1)}\leq\log_{2}(24\mathsf{w})\leq\gamma(\mathsf{w}),\qquad\forall x\geq 1.$ (5.5)
Combining (5.3), (5.5), we infer that
@@
To prove Proposition 5.2, we rely on the following two well-known lemmas. The first, Lemma 5.5 (see, for instance, [landau1965eigenvalue, landau1993density]) concerns the spectral gap in the eigenvalues $\mu_{\mathsf{w},j}$ (2.4), while the second, Lemma 5.6, pertains to the magnitude of $\varphi_{\mathsf{w},j}$ for sufficiently large index $j$. The proof of Proposition 5.2 will come after the presentation of these lemmas.
@@
Recall that the Slepian basis functions $\varphi_{\mathsf{w},j}$ serve as eigenfunctions for the operator $\mathcal{L}_{\mathsf{w}}$ (2.5), with the corresponding eigenvalues denoted by $\chi_{\mathsf{w},j}$ (2.6). Let $j\geq\frac{2\mathsf{w}}{\pi}$. Then it follows from [osipov2013prolate, Theorem 4.4] that
@@
On the one hand, when (5.8) is satisfied, we can employ [osipov2013prolate, Theorem 3.39] to obtain
@@
On the other hand, according to [bonami2014uniform, Theorem 3.1], (5.8) also yields
@@
Extend the uniform measure $\bf u$} on [1,1][-1,1] to a measure on $\mathbb{R}$ by setting ${\rm d}{\bf u}(t)=\frac{1}{2}{\rm d}t$. Under the standard energy concentration interpretation, $\mu_{\mathsf{w},j}$ represents the fraction of the total energy of $\varphi_{\mathsf{w},j}$ contained in [1,1][-1,1] (see Section 2.1). Therefore, with the normalization $\|\varphi_{\mathsf{w},j}\|_{L_{\bf u}^{2}(\mathbb{R})}=1$, we obtain
$\mu_{\mathsf{w},j}=\frac{\|\varphi_{\mathsf{w},j}\|_{L^{2}([-1,1])}}{\|\varphi_{\mathsf{w},j}\|_{L^{2}(\mathbb{R})}}=\frac{1}{\|\varphi_{\mathsf{w},j}\|_{L_{\bf u}^{2}(\mathbb{R})}}.$ (5.11)
@@
$|\varphi_{\mathsf{w},j}(x)|=\bigg|\frac{1}{2\pi}\int_{-\mathsf{w}}^{\mathsf{w}}\widehat{\varphi_{\mathsf{w},j}}(t)e^{ixt}\,dt\bigg|\leq\sqrt{\frac{\mathsf{w}}{\pi}}\|\varphi_{\mathsf{w},j}\|_{L^{2}(\mathbb{R})}=\sqrt{\frac{2\mathsf{w}}{\pi}}\|\varphi_{\mathsf{w},j}\|_{L_{\bf u}^{2}(\mathbb{R})},\qquad\forall x\in\mathbb{R}.$ (5.12)
For $j\leq\lfloor 4\mathsf{w}\rfloor-1$, applying Lemma 5.5 together with (5.11) and (5.12) yields
@@
$\alpha=\alpha(\mathcal{S}_{\Lambda},\{\vec{y}_{j}\}_{j=1}^{m})\\ :=\inf\bigg\{\frac{1}{m}\sum_{j=1}^{m}|g(y_{j})|^{2}:g\in\mathcal{S}_{\Lambda}\text{ and }\|g\|_{L_{{\bf u}}^{2}([-1,1]^{d})}=1\bigg\}.$
Then, since $\mathcal{S}_{\Lambda}$ is a finite-dimensional space, with a basis formed by the vectors $\{\varphi_{\mathsf{w},\vec{\nu}}\}_{\vec{\nu}\in\Lambda}$, we apply the Courant-Fisher min-max theorem [horn2012matrix, Theorem 4.2.6] to get
@@
where $A\in\mathbb{R}^{m\times\texttt{\#}\Lambda}$ is given by
$A_{j,\vec{\nu}}:=\dfrac{1}{\sqrt{m}}\varphi_{\mathsf{w},\vec{\nu}}(\vec{y}_{j}).$ (5.13)
Here, $\sigma_{\rm min}(A)$ denotes the minimum singular value of $A$ and $\lambda_{\rm min}(A^{*}A)$ the minimum eigenvalue of $A^{*}A$. Note that to define $A^{*}$, we assume a fixed ordering of $\vec{\nu}\in\Lambda$. A positive lower bound on $\alpha$, crucial for the existence of a unique solution to (2.8), can be further guaranteed via Monte Carlo sampling if $m$ is sufficiently large. This is formalized in the following theorem, adapted from [adcock2022sparse, Theorem 5.7].
@@
Let $\beta,\delta\in(0,1)$ and set [^3] The quantity $\kappa(\mathcal{S}_{\Lambda})$ in (5.14) is known as the $L^{\infty}$-norm of the Christoffel function of $\mathcal{S}_{\Lambda}$ [adcock2022sparse, Definition 5.4].
$\kappa(\mathcal{S}_{\Lambda}):=\max_{\vec{y}\in[-1,1]^{d}}\sum_{\vec{\nu}\in\Lambda}|\varphi_{\mathsf{w},\vec{\nu}}(\vec{y})|^{2}.$ (5.14)
If $m\geq((1-\delta)\log(1-\delta)+\delta)^{-1}\kappa(\mathcal{S}_{\Lambda})\log(\texttt{\#}\Lambda/\beta)$, then with probability at least $1-\beta$, the discrete stability constant $\alpha=\alpha(\mathcal{S}_{\Lambda},\{\vec{y}_{j}\}_{j=1}^{m})$ satisfies
@@
Evidently from (5.14), the size of $\kappa(\mathcal{S}_{\Lambda})$ is impacted by the choice of $\Lambda$. When $\Lambda$ is a hyperbolic cross in low dimensions, an upper bound for $\kappa(\mathcal{S}_{\Lambda})$ can be provided by the next proposition.